Solve by substitution. Determine whether the equation is independent, dependent or inconsistent.
y = x + 4
3y - 5x = 6
Thanks!

Answer by Mathtut(3670)

y = x + 4......eq 1
3y - 5x = 6....eq 2
:
take y's value from eq 1 and substitute it into eq 2
:
3(x+4)-5x=6
:
3x+12-5x=6
:
-2x=-6
:
x=3
:
y=x+4=3+4=7
:
this system is consistent and independent since it has one and only one solution
: